OpenStudy (awkwardpanda):

You can model the population of a certain city between 1955-2000 by the radical function

OpenStudy (shamil98):

would be helpful to know this radical function

OpenStudy (awkwardpanda):

\[P(x)=55000\sqrt{x-1945} \]

OpenStudy (awkwardpanda):

shh

OpenStudy (shamil98):

okay, you have this function, what is the question?

OpenStudy (awkwardpanda):

in which year was the population of that city 220,00*

OpenStudy (awkwardpanda):

boom

OpenStudy (anonymous):

bam and now i know why girls cant do math

OpenStudy (shamil98):

LOL

OpenStudy (shamil98):

okay, panda. first off is your radical function written correctly?

OpenStudy (anonymous):

http://radicalfunctionssp2010.weebly.com/

OpenStudy (anonymous):

go do it yourself and stop wasting my time here trolling

OpenStudy (awkwardpanda):

go awai troll

OpenStudy (anonymous):

now the panda is really awkward

OpenStudy (shamil98):

\[\huge p(x) = 55000 \sqrt{x-1945}\] ? and is that 220,000 or 22,000

OpenStudy (awkwardpanda):

220,000 .-.

OpenStudy (shamil98):

learn to type correctly -.-

OpenStudy (awkwardpanda):

no :3

OpenStudy (shamil98):

set the equation to 220,000. \[\huge 220000 = 55000\sqrt{x-1945}\] and solve for x. to get your year

OpenStudy (awkwardpanda):

thank youuu :3

OpenStudy (anonymous):

1961

OpenStudy (shamil98):

yes, 1961.

OpenStudy (anonymous):

220/55 square than add 1945

OpenStudy (shamil98):

i would rather she did the work. @trollswillrule

OpenStudy (awkwardpanda):

I'll still do the work.

OpenStudy (shamil98):

good

OpenStudy (anonymous):

omg simple equation... are u really that old as u seem on ur dp?

OpenStudy (awkwardpanda):

oh my god xD

OpenStudy (anonymous):

omg now i know why pandas are going to extinct

OpenStudy (shamil98):

LMFAOOOOO

OpenStudy (anonymous):

hey go check out michybooboo's qn

OpenStudy (anonymous):

it bothered me quite a lot lol and want someone to help me double check my answer